A rod length of 80 cm pivotted freely at its centre and placed horizontally. A 20g ping pong ball with velocity 3 m/s collides and sticks to one end of the rod causing them to rotate. The moment of inertia of ball rod system is 0.2kgm^2. What is the angular velocity of the rod?
1
Expert's answer
2019-08-08T10:57:51-0400

The law of conservation of angular momentum states

"J\omega=mvr"

So, the angular velocity of the rod

"\omega=\frac{mvr}{J}"

"=\frac{0.020\times 3\times 0.40}{0.2}=0.12\:\rm{rad/s}"